# Sample-efficient Unsupervised Policy Cloning from Ensemble Self-supervised Labeled Videos

### v1

@@ IV-D Hyper-parameter Sensitivity Analysis
In this section, we test different shift distance $s$ values in the visual shift contrast (VSC) task, as shown in Fig. 4. When $s=0$, VSC degenerated into an image discrimination task. It is also effective (much better than UPESV v/o VSC) but worse than adding a few shifts (worse than $s=1,2,4$. It means that forcing the model to ignore the global observation difference is helpful for understanding dynamics. In addition, we also try to add some time information (usually useful in RL), using two neighboring images instead of two shifted images (having the same origin) to define positive pairs for contrastive learning [48] while finding no improvements. It fits our intuition because the time contrast reduces the model’s ability to perceive neighboring distinctions, which is critical for understanding complex environmental dynamics.
TABLE II: The prediction accuracy of two video labeling models on unseen expert dataset.
@@ IV-E Robustness: Prediction Accuracy on Unseen Expert Data
Considering both BCO and our UPESV train video labeling models to infer real actions contained in expert videos, we compare the two models’ prediction accuracy on unseen expert action-labeled datasets. The results in TABLE II show the superiority of UPESV. This is consistent with their policy performance in the few-shot learning task, demonstrating the influence of additional dynamics understanding and representation learning on models’ robustness.
V Conclusion & Future Work
In this work, we propose UPESV, a novel, sample-efficient, and completely unsupervised policy learning framework to imitate policy from videos. With three organically selected and combined self-supervised tasks, UESPV models well understand both visual representation and environmental dynamics, enabling state-of-the-art few-shot policy learning in challenging procedurally-generated environments. The ability of sample-efficient video-based policy learning is a perfect fit for current challenging tasks where interactions are expensive, supervision is difficult to acquire, but video demonstrations are plentiful, e.g., robot manipulation. Crossing the gap from video games to real robots is a big challenge, which we leave in future studies.

### v2

@@ IV-D Hyper-parameter Sensitivity Analysis
In this section, we test different shift distance $s$ values in the visual shift contrast (VSC) task, as shown in Fig. 4. When $s=0$, VSC degenerated into an image discrimination task. It is also effective (much better than UPESV v/o VSC) but worse than adding a few shifts (worse than $s=1,2,4$. It means that forcing the model to ignore the global observation difference is helpful for understanding dynamics. In addition, we also try to add some time information (usually employed in RL), using two neighboring images instead of two shifted images (having the same origin) to define positive pairs for contrastive learning [51] while finding no improvements. It fits our intuition that the time contrast reduces the model’s ability to perceive neighboring distinctions, which is critical for the understanding of complex environmental dynamics.
TABLE II: Prediction accuracy comparison on unseen expert dataset.
@@ IV-E Robustness: Prediction Accuracy on Unseen Expert Data
Considering that both BCO and our UPESV train video labeling models to infer real actions contained in expert videos, we compare the two models’ prediction accuracy on unseen expert action-labeled datasets. The results in TABLE II show the superiority of UPESV. This is consistent with their performance on the policy learning tasks, demonstrating the influence of additional dynamics understanding and representation learning on models’ robustness.
V Conclusion & Limitation
In this work, we propose UPESV, a novel, sample-efficient, and unsupervised framework to learn policies from action-free videos without rewards. With three organically selected and combined self-supervised tasks, UPESV enables state-of-the-art interaction-limited policy learning in multiple challenging procedurally generated video games. There are also some limitations. Although the design of ensemble self-supervised learning improves the sample efficiency, it inevitably results in a higher computational burden due to the additional objectives. In addition, balancing multiple self-supervised tasks with different learning rates also requires additional experimental trials. Finally, there is a huge gap between video games (discrete control) and continuous control or even real robots. How to bridge this gap is unsolved in this paper, and we leave it for future study.